# Parallel Cluster-BFS and Applications to Shortest Paths

### v1

@@ 4.2   An ADO based on Landmark Labeling
Landmark labeling (LL) is one of the widely-used approaches of ADOs and probably the simplest. The basic idea is to select a subset $H$ of vertices as landmarks, and precompute the distances between each landmark $h\in H$ and all the vertices $u\in V$. When the distance between two vertices, $u$ and $v$, is queried, $\mbox{query}(u,v)$ answers the minimum $\delta(u,h)+\delta(h,v)$ over all the landmarks $h\in H$ as an estimation. We show the high-level idea about LL in Alg. 2.
@@ 5.1   Microbenchmarks for Cluster-BFS
Table 2: Tested graphs and microbenchmarks on different BFS algorithms from a cluster of vertices with size 64. The numbers endup with ‘$\times$’ are speedups, higher is better. Others are running time, lower is better. The columns “AIY”, “Ligra” in related work and “Final” show the speedup over the “Seq-BFS”. “AIY” is referred to sequential C-BFS from [3], “Ligra” is referred to parallel single BFS [45], and “Final” is referred to our parallel C-BFS. The “self-speedup” is the speedup running the algorithm in parallel over running it in sequential.
@@
We start with testing our cluster-BFS (C-BFS) as a building block. We first test the simplest case where $d=2$ and $k=w=64$, where each cluster is a star (a vertex and its 63 neighbors). Our experiment runs on 10 different clusters on each graph and the average is reported. We present a detailed experimental study for this simple case because one of our baselines, AIY, only supports sources as star-shaped clusters. Another reason is that in previous work (as well as new results in this paper), we observed that using $d=2$ gives the best overall performance for the two applications discussed in this paper. We present some studies about varying $d$ at the end of this subsection.
@@
As shown in the column “AIY”, using clusters and bit-parallelism brings up 18.8–32.0$\times$ improvement, which is uniform on different graphs. Note that here, we have $k=64$ sources, so the maximum speedup can be 64$\times$. Of course, since C-BFS is more complicated than the plain BFS, there are some constant overheads, leading to an average of 22.4$\times$ speedup in a sequential setting.
For applying thread-level parallelism, the improvement on different graphs varies greatly, from 3.91$\times$ (on smallest graphs) to up to 187$\times$ on a 96-core machine with hyperthreads. The benefit on certain graphs (e.g., OK, TW, and FT) is significant, which is over 100$\times$. One reason for the difference in improvement is the directional optimization. On some dense graphs, directional optimization may exist early in the backward traverse and thus save work (see Sec. A for more details). This is consistent with the observations in existing work [45, 8]. On average, effectively utilizing parallelism gives 27.0$\times$ speedup over plain sequential BFS.
@@
Self-relative Speedup.   In addition to the aforementioned set of baselines, we further tested C-BFS in the sequential setting to study the self-relative speedup (in column “Self-Spd.”). The speedup numbers are from 9.44$\times$ (on smallest graphs) to more than 40$\times$ (on most large graphs). In summary, the self-speedup of applying thread-level parallelism is 35.4$\times$ on average.
@@
Scalability on Different Number of Cores.   In addition to the overall self-speedup on 96 cores with hyperthreads, the scalability curve, as shown in Fig. 4, presents the self-speedup results across different core configurations. The curve demostrates that our algorithm achieves nearly linear speedup for most of the graphs, indicating efficient parallel scalability. One exception is the DBLP graph, which deviates from this pattern due to its smaller size, resulting in a running time of less than 0.02 seconds.
Influence of Cluster Diameter $d$ in C-BFS.   There are no existing implementations supporting C-BFS with general $d$. Recall that our C-BFS supports general clusters with diameter $d$ instead of star-shaped clusters (a vertex and its neighbors, where $d=2$). As shown in Thm. 3.1, the work is proportional to $d$. We tested different $d$ from $2$ to $6$ on all the graphs. We choose 10 representative graphs and show the C-BFS running time on clusters with different $d$ in Fig. 5. The full running time is shown in Tab. 4 in section 5.4. The running time increases as $d$ grows, especially when $d$ is small. A large $d$ allows for better flexibility for the shape of the cluster (the vertices can be further from each other), but significantly affects the performance. In section 5.4, we will further show that using a large $d$ in LL incurs overhead in space and time, and therefore, using $d=2$ is almost always more effective in that application.
@@ 5.3   Approximate Landmark Labeling
We now show how C-BFS can significantly improve the landmark labeling (LL) approach regarding both running time and accuracy. In general, more landmarks will lead to better accuracy for the distance queries, as the landmarks are more likely to be on or close to the shortest path between the queried vertices. Hence, by using the clusters as landmarks, we can drastically increase the number of landmarks by a factor of $w$ with $O(d)$ overhead in time and space. For simplicity, we start with the clusters with diameter $d=2$, and we will discuss the clusters for $d>2$ later. Following the optimization mentioned in section 4.2, we use each cluster as a vertex and its $w-1$ neighbors, and only store a distance and $d$ bit-subsets in the cluster distance vector.
We study the effectiveness of our approach by comparing our C-BFS-based LL with a standard solution where each landmark is one vertex. We limit the total memory usage for both algorithms to be $t$ bytes per vertex for each graph, and construct an LL-based index within this memory budget. Note that the memory usage for C-BFS requires storing the following information per cluster: the distance (1 byte), and $d$ bit-subsets ($w/8$ bytes each), while the memory usage for a regular LL is one byte (the distance) per vertex. For example, C-BFS with $w=64$ and $d=2$ needs $17$ bytes to store a cluster distance vector, and C-BFS with $w=8$ and $d=2$ only needs $3$ bytes. Thus, the memory usage depends on the word size and number of clusters we choose. For fair comparisons, we fix the memory usage per vertex in the index for different baselines. With the same memory budget, a larger $w$ results in fewer (independent) clusters, thus usually allowing for faster preprocessing time. As discussed in section 4.2, the landmarks in the same cluster are highly correlated to each other, and may not bring the same benefit as independent ones. Namely, a larger $w$ may also result in less accuracy.
@@
The experimental results suggest that, if the primary objective is to reduce preprocessing time, using C-BFS with $w=64$ will always give a more efficient version than the plain version. The precision is also superior in most cases. When achieving better precision is the top priority, C-BFS with a smaller $w$ can significantly reduce the distortion. When using $w=8$, the distortion is almost always better than both $w=64$ and the plain version, and the running time also improves the plain version on most of the graphs.
@@ 6   Related Work
Regarding the exact distance queries, Pruned Landmark Labeling (PLL) [3] is among the latest solutions for scale-free networks. Follow-up work [33] provided detailed improvements, but the lack of publicly available code prevents a direct comparison with our approach. Additionally, other techniques, such as contraction hierarchies (CH) [26] and transit nodes routing [7], focus more on large-diameter graphs like road networks.
@@ 7   Conclusion
In this paper, we present parallel implementations for cluster-BFS, which runs BFS from a cluster of vertices with diameter $d$. Our algorithm is work-efficient in theory, and also leads to high parallelism on low-diameter graphs as we tested in the experiment. We employ both bit-level and thread-level parallelism to optimize the performance. Both of them lead to significant speedup. Especially, we observed that bit-level and thread-level parallelism work well in synergy. We also show that the combination of the techniques also leads to performance improvement in two applications in distance oracles in multiple measurements of preprocessing time and accuracy, and allows our implementation to scale to much larger graphs than a sequential algorithm. Besides, our C-BFS is the first implementation that supports general clusters with diameter $d$ instead of star-shaped clusters, which give us a chance to study the benifits and overhead of choosing clusters with larger $d$ in C-BFS and in its applications.

### v2

@@ 4.2   An ADO based on Landmark Labeling
Landmark labeling (LL) is one of the widely-used approaches of ADOs and probably the simplest. The basic idea is to select a subset $H$ of vertices as landmarks, and precompute the distances between each landmark $h\in H$ and all the vertices $u\in V$. When the distance between two vertices, $u$ and $v$, is queried, $\mbox{query}(u,v)$ answers the minimum $\delta(u,h)+\delta(h,v)$ over all the landmarks $h\in H$ as an estimation. We show the high-level idea of LL in Alg. 2.
@@ 5.1   Microbenchmarks for Cluster-BFS
Table 2: Tested graphs and microbenchmarks on different BFS algorithms from a cluster of vertices with size 64. The numbers followed by ‘$\times$’ are speedups, higher is better. Others are running time, lower is better. The columns “AIY”, “Ligra” in related work and “Final” show the speedup over the “Seq-BFS”. “AIY” is referred to sequential C-BFS from [3], “Ligra” is referred to parallel single BFS [45], and “Final” is referred to our parallel C-BFS. The “self-speedup” is the speedup running the algorithm in parallel over running it in sequential.
@@
We start with testing our cluster-BFS (C-BFS) as a building block. We first test the simplest case where $d=2$ and $k=w=64$, where each cluster is a star (a vertex and its up to 63 neighbors). We present a detailed experimental study for this simple case because one of our baselines, AIY, only supports sources as star-shaped clusters. Another reason is that in previous work (as well as new results in this paper), we observed that using $d=2$ gives the best overall performance for the two applications discussed in this paper. We present some studies about varying $d$ at the end of this subsection.
@@
As shown in the column “AIY”, using clusters and bit-parallelism gets up to 18.8–32.0$\times$ improvement, which is uniform on different graphs. Note that here, we have $k=64$ sources, so the maximum speedup can be 64$\times$. Since C-BFS is more complicated than the plain BFS, there are some constant overheads, resulting in an average 22.4$\times$ speedup in a sequential setting.
For applying thread-level parallelism, the improvement on different graphs varies greatly, from 3.91$\times$ (on smallest graphs) to up to 187$\times$ on a 96-core machine with hyperthreads. The benefit on certain graphs (e.g., OK, TW, and FT) is significant, which is over 100$\times$. One reason for the difference in improvement is the directional optimization. On some dense graphs, the backward step can be applied across many of the rounds and thus save significant work (see Sec. A for more details). Another reason is that there is more parallelism available on larger graphs. This is consistent with the observations in prior work [45, 8]. On average, effectively utilizing parallelism gives 27.0$\times$ speedup over plain sequential BFS.
@@
Self-relative Speedup and Scalability.   In addition to the aforementioned set of baselines, we further tested C-BFS in the sequential setting to study the self-relative speedup (in column “Self-Spd.”). The speedup numbers are from 9.44$\times$ (on smallest graphs) to more than 40$\times$ (on most large graphs). In summary, the self-speedup of applying thread-level parallelism is 35.4$\times$ on average.
@@
In addition to the overall self-relative speedup on 96 cores with hyperthreads, the scalability curve, as shown in Fig. 4, presents the self-relative speedup results across different number of cores. The curve demonstrates that our algorithm achieves nearly linear speedup for most of the graphs, indicating efficient parallel scalability. One exception is the DBLP graph, which deviates from this pattern due to its smaller size.
Influence of Cluster Diameter $d$ in C-BFS.   There are no existing implementations supporting C-BFS with general $d$. Recall that our C-BFS supports general clusters with diameter $d$ instead of star-shaped clusters (a vertex and its neighbors, where $d=2$). As shown in Thm. 3.1, the work is proportional to $d$. We tested different $d$ from $2$ to $6$ on all the graphs. We choose 10 representative graphs and show the C-BFS running time on clusters with different $d$ in Fig. 5. The full running time is shown in Tab. 4 in section 5.4. The running time increases as $d$ grows. A large $d$ allows for better flexibility for the shape of the cluster (the vertices can be further from each other), but significantly affects the performance. In section 5.4, we will further show that using a large $d$ in LL incurs overhead in space and time, and therefore, using $d=2$ is almost always more effective in applications.
@@ 5.3   Approximate Landmark Labeling
We now show how C-BFS can significantly improve the landmark labeling (LL) approach with respect to both running time and accuracy. In general, more landmarks will lead to better accuracy for the distance queries, as the landmarks are more likely to be on or close to the shortest path between the queried vertices. Hence, by using the clusters as landmarks, we can drastically increase the number of landmarks by a factor of $w$ with $O(d)$ overhead in time and space. For simplicity, we start by considering clusters with diameter $d=2$, and later discuss clusters with $d>2$. Following the optimization mentioned in section 4.2, we use each cluster as a vertex and its $w-1$ neighbors, and only store a distance and $d$ bit-subsets in the cluster distance vector.
We study the effectiveness of our approach by comparing our C-BFS-based LL with a standard solution where each landmark is one vertex. We limit the total memory usage for both algorithms to with a parameter of $t$ bytes per vertex for each graph, and construct an LL-based index within this memory budget. For C-BFS, memory usage per cluster includes the distance (1 byte) and $d$ bit-subsets ($w/8$ bytes each) per vertex, while the memory usage for a regular LL is one byte (the distance) per vertex. For example, C-BFS with $w=64$ and $d=2$ needs $17$ bytes to store a cluster distance vector, and C-BFS with $w=8$ and $d=2$ only needs $3$ bytes. Thus, the memory usage depends on the word size and number of clusters we choose. For fair comparisons, we fix the memory usage per vertex in the index for different baselines. With the same memory budget, a larger $w$ results in fewer (independent) clusters, typically allowing faster preprocessing. As discussed in section 4.2, the landmarks in the same cluster are highly correlated to each other, and may not bring the same benefit as independent ones. Thus, a larger $w$ may also result in less accuracy.
@@
The experimental results suggest that, if the primary objective is to reduce preprocessing time, using C-BFS with $w=64$ will always give a more efficient version than the plain version. The precision is also superior in most cases. When achieving better precision is the top priority, C-BFS with a smaller $w$ can significantly reduce the distortion. When using $w=8$, the distortion is almost always better than both $w=64$ and the plain version, and the running time also improves over the plain version on most of the graphs.
@@ 6   Related Work
Regarding the exact distance queries, Pruned Landmark Labeling (PLL) [3] is among the latest solutions for scale-free networks. Li et al. [33] showed another implementation, but they did not release their code, so we cannot compare their running time with ours. Other techniques, such as contraction hierarchies (CH) [26] and transit nodes routing [7], focus more on large-diameter graphs like road networks.
@@ 7   Conclusion
In this paper, we present parallel implementations for cluster-BFS, which runs BFS from a cluster of vertices with diameter $d$. Our algorithm is work-efficient in theory, and also leads to high parallelism on low-diameter graphs as we tested in the experiments. We employ both bit-level and thread-level parallelism to optimize the performance. Both of them lead to significant speedup. Especially, we observed that bit-level and thread-level parallelism work well in synergy. We also show that the combination of the techniques also leads to performance improvement in two applications in distance oracles in multiple measurements of preprocessing time and accuracy, and allows our implementation to scale to much larger graphs than a sequential algorithm. Besides, our C-BFS is the first implementation that supports general clusters with diameter $d$ instead of star-shaped clusters, which give us a chance to study the benifits and overhead of choosing clusters with larger $d$ in C-BFS and in its applications.
Acknowledgments
This work is supported by NSF grants CCF-1919223, CCF-2103483, CCF-2119352, IIS-2227669, TI-2346223, and NSF CAREER Awards CCF-2238358 and CCF-2339310, the UCR Regents Faculty Development Award, and the Google Research Scholar Program.